OpenStudy (ineedhelpnow):

3x/5-x=x/10-5/2 solve

OpenStudy (anonymous):

(3x)/(5)-x=(x)/(10)-(5)/(2) Since (x)/(10) contains the variable to solve for, move it to the left-hand side of the equation by subtracting (x)/(10) from both sides. (3x)/(5)-x-(x)/(10)=-(5)/(2) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 10. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. -x*(10)/(10)-(x)/(10)+(3x)/(5)*(2)/(2)=-(5)/(2) Complete the multiplication to produce a denominator of 10 in each expression. -(10x)/(10)-(x)/(10)+(6x)/(10)=-(5)/(2) Combine the numerators of all expressions that have common denominators. (-10x-x+6x)/(10)=-(5)/(2) Combine all like terms in the numerator. (-5x)/(10)=-(5)/(2) Move the minus sign from the numerator to the front of the expression. -(5x)/(10)=-(5)/(2) Reduce the expression -(5x)/(10) by removing a factor of 5 from the numerator and denominator. -(x)/(2)=-(5)/(2) Multiply each term in the equation by 2. -(x)/(2)*2=-(5)/(2)*2 Simplify the left-hand side of the equation by canceling the common terms. -x=-(5)/(2)*2 Simplify the right-hand side of the equation by simplifying each term. -x=-5 Multiply each term in the equation by -1. -x*-1=-5*-1 Multiply -x by -1 to get x. x=-5*-1 Multiply -5 by -1 to get 5. x=5

OpenStudy (anonymous):

^

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

x=5